please help me in this age word problem : a woman is now 40 years old and her daughter is 14 years old. in how many years will the mother be exactly twice as old as her daughter?
Answer by nerdybill(7384)

Let x = number of years when mother will be twice as old as daughter
then
"mother's age" = twice "daughter's age"
40+x = 2(14+x)
40+x = 28+2x
40 = 28+x
12 = x
.
In 12 years, Mom will be twice daughter's age